OpenStudy (anonymous):

What is the solution to the system of equations? x+y=5 x-2y=2 ( , )

Vocaloid (vocaloid):

hint: start by subtracting the two equations

Vocaloid (vocaloid):

what is x - x ?

OpenStudy (anonymous):

Wait how would I do that

Vocaloid (vocaloid):

start with x take away x what's left?

OpenStudy (anonymous):

nothing

Vocaloid (vocaloid):

good, so x - x = 0

Vocaloid (vocaloid):

now what is y - (-2y)?

OpenStudy (anonymous):

so I would have to do the parentheses first?

OpenStudy (anonymous):

How I still have to figure out Y

Vocaloid (vocaloid):

you don't, you just leave y as it is

Vocaloid (vocaloid):

hint: 1-(-2) = 1 + 2 y-(-2y) = ______ fill in the blank

OpenStudy (anonymous):

It's confusing

Vocaloid (vocaloid):

y-(-2y) = y + 2y = ?

Vocaloid (vocaloid):

I start with y then I add two more y's how many y's do I have?

OpenStudy (anonymous):

3

Vocaloid (vocaloid):

good so y + 2y = 3y

OpenStudy (anonymous):

ohh lol I suck at math

Vocaloid (vocaloid):

then we finish subtracting 5 - 2 = 3 that gives us 3y = 3 so what does y equal?

OpenStudy (anonymous):

1

Vocaloid (vocaloid):

good now, we know that y = 1 so if x + y = 5, what does x equal?

OpenStudy (anonymous):

4

Vocaloid (vocaloid):

good so our answer is (4,1)

OpenStudy (anonymous):

Omg Thank You so much!!